OpenStudy (anonymous):

solve for x 3log_(4)x+log_(4)2-log_(4)(x-2) show steps if you can!

OpenStudy (anonymous):

Can you please put it in the equation. I can't understand what it says.

OpenStudy (anonymous):

*equation?

OpenStudy (mathstudent55):

\(\Large 3 \log_4 x+ \log_4 2- \log_4 (x-2)\)

OpenStudy (anonymous):

\[3\log_4x+\log_42-\log_4(x-2)\]

OpenStudy (mathstudent55):

Is that what you meant? There is no equal sign, so it is not an equation. What do you need to do with it? Write it as a single log? Or is the equal sign missing?

OpenStudy (anonymous):

the question just said "Simplify."

OpenStudy (anonymous):

I think I just need to write it as a single log since there is no equal sign.

OpenStudy (mathstudent55):

Ok. I think what that means is to write the entire thing as a single log.

OpenStudy (mathstudent55):

You need to know these three rules of logs: \(\log a + \log b = \log ab\) \(\log a - \log b = \log \dfrac{a}{b} \) \(\log a^n = n \log a\) The third rule is not needed in this problem, but you need to know it.

OpenStudy (anonymous):

yeah, i have those in my notes.

OpenStudy (mathstudent55):

Look only at the part in red below for now. Use the first rule. How do you combine that sum of logs into one single log? \(\Large \color{red}{3 \log_4 x+ \log_4 2}- \log_4 (x-2)\)

OpenStudy (mathstudent55):

BTW, I'm sorry, but you do need the third rule. Use the rule of the exponent first on the first log. Then combine the two red logs into one using the first rule.

OpenStudy (mathstudent55):

Use the third rule (exponent) on the green log below. \(\Large \color{green}{3 \log_4 x} + \log_4 2- \log_4 (x-2)\)

OpenStudy (mathstudent55):

Can you complete the bottom equation below?